solve for x: 2+3x=-1
Answer by vheroli(126)

2+3x = -1 :First, leave the x on one side by transposing +2 on the other side
3x = -1-2 :Then simplify
3x = -3 :Next is divide both sides by 3 to leave x alone.
3x/3 = -3/3 :Then simplify
x=-1
Answer: x=-1